OpenStudy (anonymous):

Solve for x and simplify your answer: 3x^2+3=4x

OpenStudy (anonymous):

\[3x^2-4x+3=0\] now factor or find discriminant

OpenStudy (anonymous):

x=2+i(5)/3 or x=2-i(5)/3 is that correct?

OpenStudy (anonymous):

(5) is the radical

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (anonymous):

yeah it's correct